hello. i have a problem that i do not understand,can you help me. It says:
In Tess'store, four pounds of pears cost $6. In Dorothy's store, 10 pounds of pears are sold for $15. Bubble in 1 if it costs more, 2 if it costs less, and 3 if it is the same to buy pears in Dorothy's store.

Found 2 solutions by Alan3354, stanbon:
Answer by Alan3354(69443)

hello. i have a problem that i do not understand,can you help me. It says:
In Tess'store, four pounds of pears cost $6. In Dorothy's store, 10 pounds of pears are sold for $15. Bubble in 1 if it costs more, 2 if it costs less, and 3 if it is the same to buy pears in Dorothy's store.
--------------
In Tess's store, they cost $1.50/pound ($6/4)
In Dorothy's store, they're $1.50/pound ($15/10)

Answer by stanbon(75887)

In Tess'store, four pounds of pears cost $6. In Dorothy's store, 10 pounds of pears are sold for $15. Bubble in 1 if it costs more, 2 if it costs less, and 3 if it is the same to buy pears in Dorothy's store.
------------
Tess's store charges 6/4 = 3/2 dollars per pound
---
Dorothy's store charges 15/10 = 3/2 dollars per pound
-----
So, Bubble in "3".
Cheers,
Stan H.
===========